myininaya (myininaya):

\[\text{ Evaluate } \int \text{ HappyHolidays } d(\text{ HappyHolidays }) \]

Nnesha (nnesha):

lol

OpenStudy (anonymous):

it will be some thing with C (celebrate )

OpenStudy (anonymous):

\[\frac{\text{Happy Holidays}^2}{2}+Chirstmas\]

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

\[\frac{\text{Happy Holidays}^2}{2}+Christmas\]

Nnesha (nnesha):

\(\large\color{green}{ {\rm holidays^2}~~~+ winter^2 = happy^2 }\) :

Nnesha (nnesha):

lol

OpenStudy (anonymous):

Also it reminds me of, solve \[\huge y = \frac{ \ln(\frac{ x }{ m }-sa) }{ r^2 }\]

OpenStudy (anonymous):

why Christmas = Xmas ?

myininaya (myininaya):

 \[\cos(\theta)=\frac{ \text{ winter }}{ \text{ happy } }\]

Nnesha (nnesha):